Maddox takes off on his jet ski so they have a combined mass of 350.51kg. The skis have a thrust of 1,256.46N and a coefficient of kinetic friction on snow of 0.07. Maddox failed to check the gas tank and ran out of gas after only 19.72 s. What is Maddox's top speed?
Answer #1

Frictional force, Ff = umg = 0.07 x 350.51 x 9.8 = 240.45 N

Net force, F = 1256.46 - 240.45 = 1016.01 N

Acceleration, a = F/m = 1016.01/350.51 = 2.90 m/s^2

Velocity, v = u + at = 0 + 2.90 x 19.72 = 57.16 m/s

Comment in case any doubt please rate my answer....